OpenStudy (anonymous):

Can somebody help me figure out the limit as x approaches infinity from the left of (7/x - x/2)?

OpenStudy (anonymous):

combine the two fractions

OpenStudy (anonymous):

Did that and got DNE.

OpenStudy (anonymous):

well, DNE is technically correct, but some people consider the case +/- infinity to be the limit of the function

OpenStudy (anonymous):

Well, webassign won't take either of those answers. I have tried them both.

OpenStudy (anonymous):

did you try *negative* infinity?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

my question is, what does "approaches infinity from the left" mean?

OpenStudy (jdoe0001):

hmm right

OpenStudy (anonymous):

you can go to \(\infty\) only from the left

OpenStudy (anonymous):

so maybe there is a typo in the question

OpenStudy (anonymous):

Okay, it wanted infinity... that seems like half an answer to me.

OpenStudy (jdoe0001):

I gather he meant \(\Large \bf lim_{x\to + \infty^-}\)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

my guess is it is \[\lim_{x\to -\infty}\]

OpenStudy (anonymous):

in which case \(\infty\) would be the right answer, assuming you take that as a limit

OpenStudy (anonymous):

Yeah, you are right. I hate limits. Can you explain, using rules of infinite limits why, DNE wouldn't be the answer?

OpenStudy (anonymous):

it would be, unless you admit \(\infty\) or \(-\infty\) as an answer it is really talking about "end behavior"

OpenStudy (anonymous):

I'm guessing DNE is preserved for the limit of a function of at a single point.

OpenStudy (anonymous):

So, it being a one sided limit is the issue. If it were just the limit as x approaches infinity from both sides, it wouldn't exist.

OpenStudy (jdoe0001):

@jcates00 because you're approaching a value that's never defined, since -infinity would be by definition just that, thus it's \(\bf DNE \iff -\infty\)